The threshold wavelength of tungsten is 2300 angstroms. If ultra violet light of wavelength 1800 angstroms is incident on it,then what would be the maximum kinetic energy of photoelectrons?
 
 

Grade:12

1 Answers

Arun
5 years ago
wavelength(λ) of UV light= 188 nm 
Energy of incident rays = hc/λ (h=planks constant,c=speed of light) =1242 eV-nm/188 nm=6.606 eV 
Threshold λ=230 nm 
Threshold energy=work function ( Φ )=1242eV-nm/230 nm= 5.4 eV. 
By einstein photo-electric effect, 

Energy Incident=Φ + Maximum K.E. 

Thus maximum kinetic energy of emitted electrons would be=6.606 - 5.4 = 1.206 eV.
 
RegaRds
Arun